Please help me solve this equation: 15-3x=12-2x

Thank-you so much! I am revising for a test and preferably need it to be done step by step!

Answer by Scorinitron(19)

____________________________________________________
Ok I'll explain everything I did step by step
Here's your Problem
15-3x=12-2x


Since -2x contains the variable to solve for, move it to the left-hand side of the equation by adding 2x to both sides.
-3x+15+2x=12


Since -3x and 2x are like terms, subtract 2x from -3x to get -x.
-x+15=12


Since 15 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 15 from both sides.
-x=-15+12


Add 12 to -15 to get -3.
-x=-3


Multiply each term in the equation by -1.


-x*-1=-3*-1

Multiply -x by -1 to get x.

x=-3*-1


Multiply -3 by -1 to get 3.


x=3
Your Answer is 3
NOW CHECK YOUR ANSWER:
__________________________

Now that you have your answer you have to check your problem to see if it's correct. You accomplish this by replacing the variable x with what you got which was 3 so go back to your original problem


15-3x=12-2x


and you replace x with 3 since we figured out x=3

so

15-3*(3)=12-2*(3)


Multiply -3 by each term inside the parentheses.


15-9=12-2*(3)


Subtract 9 from 15 to get 6.


6=12-2*(3)


Multiply -2 by each term inside the parentheses.


6=12-6


Subtract 6 from 12 to get 6.
6=6
Since 6=6, the equation will always be true.
So x=3 which means were correct!